Circular No. 6227

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 73P/SCHWASSMANN-WACHMANN 3 J. Crovisier, N. Biver, D. Bockelée-Morvan, P. Colom, E. Gérard, L.

Jorda, and H. Rauer, Observatoire de Paris-Meudon, observing with the Nançay Radio Telescope, report that observations of OH lines reveal an outburst of this comet, with average Q(OH) = (1.2 ± 0.12) × 10^{29} s^{−1} during Sept. 8–12, compared with a 3σ upper limit (average) of 4 × 10^{28} s^{−1} during Sept. 1–5. Observations continuing on Sept. 13 suggest that the OH lines are even stronger than on Sept. 12.

SUPERNOVA 1995Y IN NGC 410 J. Jahn, Bodenteich, Germany, reports that E. Jung (Marl, Germany)

measured R = 15.5 on Sept. 4.947 UT. Jung also provides astrometry consistent with the values reported on IAUC 6222, his position end figures being 59^{s}.53, 42′′.2 (equinox 2000.0; GSC stars).

NOVA CASSIOPEIAE 1995 Additional prediscovery magnitudes from K. Kanatsu, Yonago, Tot-

tori, Japan (T-Max 400 film; GSC comparison stars): July 31.585 UT, 9.8; Aug. 18.596, 9.2; 22.593, 9.2. Further visual magnitude estimates (cf. IAUC 6214): Aug. 29.86 UT, 9.5 (G. Badin, Pordenone, Italy); 30.85, 9.0 (M. V. Zanotta, Como, Italy); Sept. 3.882, 8.9 (E. Broens, Mol, Belgium); 4.80, 9.0 (L. Kiss, Szeged, Hungary); 6.908, 8.8 (S. Meister, Buelach, Switzerland); 10.83, 8.9 (Kiss); 13.20, 8.9 (C. E. Spratt, Victoria, BC).

COMET C/1995 O1 (HALE-BOPP) Further total visual (and CCD) magnitude estimates: Aug. 24.91 UT,

10.3 (J. G. de S. Aguiar, Campinas, São Paulo, Brazil, 11×80 binoculars); 26.19, 9.8 (C. S. Morris, Lockwood Valley, CA, 0.26-m reflector); 26.89, 9.7 (J. Carvajal, Avila, Spain, 0.44-m reflector); 29.22, 10.7 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 30.44, 10.0 (D. A. J. Seargent, The Entrance, N.S.W., 0.25-m reflector); 31.21, 10.0 (R. Keen, Mt. Thorodin, CO, 0.15-m reflector); Sept. 1.46, 10.1 (A. Nakamura, Kuma, Ehime, Japan, 0.60-m reflector + CCD); 2.17, 10.2 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 12.12, 10.6 (Hale).

1995 September 13 Daniel W. E. Green